A consumer has $25 in budget to purchase goods X & Y. Assume Px = 3 and Py=1. What is the maximum amount of X that she can buy?
1
Expert's answer
2020-02-17T08:46:19-0500

Amount=$25

Px= 3

Py= 1

budget line,"bl=Px\/Py"

bl= 3/1= 3

The amount of good X that can be bought is:

25/3= 8units